# Formal Definition and Implementation of Reproducibility Tenets for Computational Workflows

### v1

@@ Abstract
Computational workflow management systems power contemporary data-intensive sciences. The slowly resolving reproducibility crisis presents both a sobering warning and an opportunity to iterate on what science and data processing entails. The Square Kilometre Array (SKA), the world’s largest radio telescope, is among the most extensive scientific projects underway and presents grand scientific collaboration and data-processing challenges. This work presents a scale and system-agnostic computational workflow model and extends five well-known reproducibility tenets into seven defined for our workflow model. Subsequent implementation of these definitions, powered by blockchain primitives, into the Data Activated Flow Graph Engine (DALiuGE), a workflow management system for the SKA, demonstrates the possibility of facilitating automatic formal verification of scientific quality in amortized constant time. We validate our approach with a simple yet representative astronomical processing task; filtering a noisy signal with a lowpass filter with both CPU and GPU methods. Our framework illuminates otherwise obscure scientific discrepancies and similarities between principally identical workflow executions.
@@ 1 Introduction
Reuse is the ultimate goal of science; by building trust in others’ observations, theories, models, and experiences, humanity explores the universe. The ability to reproduce elements of scientific works increases our trust in their findings. The reproducibility crisis [1] heightened scrutiny and discussion around what highly trusted modern science entails. The SKA is an already storied behemoth of academic collaboration that demands paying attention to scientific reproducibility. Radio astronomy relies heavily on computational methods but is fundamentally an experimental, observational science, and hence, it presents a curious case for scientific reproducibility. Seeking a balance between the world of formal software engineering and the exploratory nature of scientific investigation has proved challenging to achieve [2, 3]. The SKA, representing an order of magnitude increase in data-processing requirements on current observatories [4], benefits from a focus on building scientific reproducibility into its instrumentation, providing an opportunity to improve what reproducibility practically means for astronomy and other computationally powered sciences. The literature widely debates the precise meaning of scientific reproducibility and replicability [5, 6] and, concerning computing, focuses on ‘forwards facing’ approaches to guarantee the future reproducibility of current computations. While practically valuable, the general trend toward adhering to specific technologies and conventions is potentially ruinous from a philosophical standpoint; how can we know that our efforts to make scientific works reproducible succeed? This work presents and validates a scale and system agnostic framework for ‘backwards facing’ reproducibility tests on computational workflow executions, facilitating efficient verification of scientific reproducibility for astronomical workflows. Specifically, we make the following contributions:
@@
    Present a scale-agnostic computational workflow model suitable for computational reproducibility discussions.
@@
    Describe a hash-graph-based workflow signature method, termed BlockDAGs, based on the scale-agnostic model allowing for constant time assessment of our reproducibility tenets.
@@
    Implement this signature mechanism in the DALiuGE [7] workflow management system.
@@
    Demonstrate the efficacy of our model, workflow signature method, and reproducibility tenets with a demonstrative lowpass filter workflow. Our method uncovers obscure differences between in-principal identical workflows.
@@ 2 Related Work
Reproducibility is fundamental to science, and computational reproducibility in support of quality science is a long-standing goal in the field, but one where the very concept is confused. Computational workflow managers are among the most powerful tools available in this quest, which, alongside contemporary software engineering techniques and provenance capture, are moving towards a world of fully rerunnable, repeatable, reproducible, and replicable research. Claerbout [8] coined the concept of computational reproducibility in scientific computing, and Peng [9] elaborates on what is needed to achieve computational reproducibility in science. Drummond [10] separates the concepts of reproducibility and applicability. Barba [5] reviews the myriad of terminologies used to describe computational reproducibility as Benureau and Rougier [11] provide a framework separation rerunning, repeating, reproducing, and replicating computational works. Central to these discussions is a separation between bitwise precise replication of exact results and conceptual reproductions of an existing work independent of precise implementation. Our definitions of these reoproducibility tenets unify these two views. Current suggested best practice in pursuit of scientific reproducibility range simple rules [12], adapting modern tooling from industry including notebook environments [13, 14, 15], sharing code [16] and publishing openly [17, 18]. Related to the signature approach we present in this work, Bellini investigates the feasibility of building a DOI-like dataset identifier leveraging public blockchains [19].
@@
Highly scalable and trustable computing is critical to conquering our largest scientific endeavors [26], and as we move towards reproducible computational workflow management, we must also ensure our efforts function as expected; a key motivation behind our signature-based approach to workflow reproducibility.
@@ 3 Computational Workflows
Computational workflows (herein, ‘workflows’) are structured assemblies of computational tasks and data stores [27] and underpin some of the world’s most extensive science projects, including the SKA [4]. Workflows separate domain-specific software knowledge from scientific methodology, making increasingly complex computations possible for increasingly more scientists, and have seen broad uptake in the past three decades [28]. Many workflow management systems provide several flavors of workflows for specific science domains [29], and thus, any attempt to address workflow reproducibility should be tool-agnostic. DALiuGE [7], [30] is a workflow management system developed to execute workflows at the scale the SKA requires. The unique data-driven design, decentralized execution, and a careful share-nothing architecture enable this. Moreover, DALiuGE strictly separates logical workflow design from physical workflow execution and thus allows astronomers to design workflows without intimate component knowledge and, more importantly, facilitates the broad reuse of logical workflows. Derived from the DALiuGE design, our scale and system-agnostic generalized workflow model alternates data artifacts with computational tasks and separates logical elements from physically executed elements. This separation provides more flexibility in reasoning about reproducing a workflow execution.
@@ 3.1 Scale and System Agnostic Workflow Model
Figure 1 presents the generalized Unified Modeling Language (UML) workflow model to make our reasoning about testing workflow executions concrete and applicable to arbitrary workflow management systems.
Figure 1: A UML diagram depicting our arbitrary-scale workflow definition.
@@
A logical task describes a task characterized by details like programming language, algorithm, and required options and parameters.
@@
A UML definition makes all modeled workflows well-defined software in their own right and permits native workflow implementation independent of using a workflow management system. Moreover, we separate logical and physical workflow components. This separation means we can reason and compare logical workflow designs irrespective of their eventual implementation in software as physical components. Moreover, several physical components can implement a single logical workflow component, as is the case in large-scale workflows. Finally, our model’s requirement to alternate computing tasks and data artifacts ensures that a workflow description captures all interim and transient data artifacts. We note that this model applies only to conventional task-based workflow management systems. While further model formalization, including ontologies for generalized scientific processes, allows accurate descriptions of data provenance [31], further formalization is beyond the scope of this research, whose focus lies on direct software implementation and testing of high-level reproducibility concerns.
3.1.1 Mapping the generalized workflow model to DALiuGE
@@
For this study, we have implemented the generalized ideas described above in DALiuGE since we have complete code control and in-depth knowledge of the code structure. DALiuGE specifies workflows as graphs of components and separates workflows into two main layers, logical and physical, and then sub-divides each of those into template and non-template versions. Workflows are developed as logical templates comprised of exclusively logical components and control structures. Scientists will design a Logical Graph Template (LGT) several months or years before the actual processing and for extensive reuse. LGTs expose configuration options to make them more specific for concrete execution. A Logical Graph (LG) is a LGT with a concrete configuration. LG specification occurs on the order of weeks before execution. Isolating details of storage components only yields a Physical Graph Stencil (PGS), which is defined only for the purpose of provenance storage and is technically a sub-graph of a LG. DALiuGE then unrolls the LG into Physical Graph Template (PGT), resolving control structures and presenting the net computation required. This process is unique for each LG, but the result can vary depending on the translator algorithm used. A physical graph template is a directed acyclic graph (DAG), a common representation for computational workflows [23]. During the scheduling process, DALiuGE partitions a PGT into sub-graphs based on available compute resources, resulting in a Physical Graph (PG). This step is performed just before execution since it maps the physical workflow components to the available resources. Each component of the DAG is a fully specified computational task or data artifact. Finally, the DALiuGE daemon triggers the first physical components in a graph, but after, components communicate with each other in a fully decentralized manner. Components can store runtime-specific information such as log files or data summaries, which together form a Runtime Graph (RG) after graph execution has been completed.
Table 1: DALiuGE to workflow model element mapping.
Workflow Model Element DALiuGE Element
Component Component
Logical Task Logical Component
Physical Task Application Component
Logical Data Artifact Logical Data Component
Physical Data Artifact Data Component
Logical Workflow Logical Graph
Physical Workflow Physical Graph
Table 1 summarizes the mapping of our generic workflow model to DALiuGE components. Adding this layer of abstraction permits multiple workflow system mappings, reproducibility definitions, and testing mechanisms to apply to any mapped system.
@@ 4 Reproducible Workflows
Contemporary science is complex; variability in the design and execution of complex data analysis pipelines can be scientifically significant [32]. Most attempts to ensure workflows are reproducible involve analyzing artifact provenance, describing precisely how pieces of data came to exist [31]. We build on this lineage, but instead of designing provenance capture to describe the entire technical process, we formulate reproducibility tests, singling out parts of workflow executions expected to be kept constant between executions. Doing so allows for flexible workflow characterization and does not limit our approach to a specific workflow management system, a critical conceptual requirement for the multi-decade-spanning SKA project and the broader scientific field. Exploiting the structured nature of workflows provides constant time provenance comparison, suitably efficient for deployment at extreme data scales. Our example workflow investigates common difficulties encountered when reproducing computational workflows. A lowpass filter workflow implemented with various methods focuses on a single component and, in the case of Fast Fourier Transform (FFT) implementations, utilizes an operation critical to many astronomical workflows. Our selected example workflows are indicative of common difficulties encountered when reproducing computational workflows.
@@ 4.1 Reproducibility Tenets
Melding computing into satisfyingly reproducible science is a well-discussed and unsolved problem [1, 33, 6]. Current efforts range from asserting standards on individual pieces of code [11], enacting well-defined data management protocols [22, 34], containerizing whole workflows [35], measures on behalf of journal publishers [36], and conference organizers [37] and careful provenance production and collection [31]. While certainly practical, we propose that forward-facing measures insufficiently integrate computing into science; they are ultimately empirical and require an inductive step to evaluate their efficacy. Our proposed approach demarcates implementation from abstract computational ideas, data and method reproduction, workflow similarities between workflow management systems, and any combination of these traits. We present seven reproducibility tenets formulated as workflow component invariants, each defining a testable assertion rather than a series of prescribed rules. For each tenet, we provide an unambiguous definition, justification, and a few motivating example scenarios, providing practical motivation behind our definition.
@@
We use the popular terminology to rerun, repeat, recompute, reproduce, and replicate computational workflows as these terms capture intuitive concepts around scientific reproducibility [38, 11, 5, 39]. Table 2 summarizes our definitions, and the later section 5.2 describes the precise provenance information the chosen workflow system has to include to satisfy each tenet.

### v2

@@ Abstract
Computational workflow management systems power contemporary data-intensive sciences. The slowly resolving reproducibility crisis presents both a sobering warning and an opportunity to iterate on what science and data processing entails. The Square Kilometre Array (SKA), the world’s largest radio telescope, is among the most extensive scientific projects underway and presents grand scientific collaboration and data-processing challenges. In this work, we aim to improve the ability of workflow management systems to facilitate reproducible, high-quality science. This work presents a scale and system-agnostic computational workflow model and extends five well-known reproducibility concepts into seven well-defined tenets for this workflow model. Additionally, we present a method to construct workflow execution signatures using cryptographic primitives in amortized constant time. We combine these three concepts and provide a concrete implementation in Data Activated Flow Graph Engine (DALiuGE), a workflow management system for the SKA to embed specific provenance information into workflow signatures, demonstrating the possibility of facilitating automatic formal verification of scientific quality in amortized constant time. We validate our approach with a simple yet representative astronomical processing task: filtering a noisy signal with a lowpass filter using CPU and GPU methods. This example shows the practicality and efficacy of combining formal tenet definitions with a workflow signature generation mechanism. Our framework, spanning formal UML specification, principled provenance information collection based on reproducibility tenets, and finally, a concrete example implementation in DALiuGE illuminates otherwise obscure scientific discrepancies and similarities between principally identical workflow executions.
@@ 1 Introduction
Reuse is the ultimate goal of science; by building trust in observations, theories, models, and experiences, humanity explores the universe. The ability to reproduce elements of scientific works increases our trust in their findings. The reproducibility crisis [1] heightened scrutiny and discussion around what highly trusted modern science entails. The SKA is an already storied behemoth of academic collaboration that demands paying attention to scientific reproducibility. Radio astronomy relies heavily on computational methods but is fundamentally an experimental, observational science, and hence, it presents a curious case for scientific reproducibility. Seeking a balance between the world of formal software engineering and the exploratory nature of scientific investigation has already proved challenging to achieve [2, 3]. The SKA, representing an order of magnitude increase in data-processing requirements on current observatories [4], benefits from a focus on building scientific reproducibility into its instrumentation, providing an opportunity to improve what reproducibility practically means for astronomy and other computationally powered sciences. The literature widely debates the precise meaning of scientific reproducibility and replicability [5, 6] and, concerning computing, focuses on ‘forwards facing’ approaches to guarantee the future reproducibility of current computations. While practically valuable, the general trend toward adhering to specific technologies and conventions is potentially ruinous from a philosophical standpoint; how can we know that our efforts to make scientific works reproducible succeed? This work presents and validates a scale and system agnostic framework for ‘backwards facing’ reproducibility tests on computational workflow executions, facilitating efficient verification of scientific reproducibility for astronomical workflows. Specifically, we make the following contributions:
@@
    Present a scale-agnostic computational workflow UML model suitable for computational reproducibility discussions.
@@
    Describe a hash-graph-based workflow signature method, termed BlockDAGs, based on the scale-agnostic model allowing for constant time generation of workflow signatures.
@@
    Implement this signature mechanism in the DALiuGE [7] workflow management system combined with provenance information collection guided by the workflow model and tenet definitions.
@@
    Demonstrate the efficacy of the workflow model, signature method, and reproducibility tenet definitions with a demonstrative lowpass filter workflow.
While the reproducibility tenets we present have appeared in related contexts before, the main contribution of our work lies in integrating these tenets together into a novel testing mechanism. This mechanism conceptually and technically synthesizes elements from across the literature, offering a unified approach that advances beyond a simple presentation of existing tenets or by taking a singular view of computational workflow reproducibility.
@@ 2 Related Work
Reproducibility is fundamental to science, and computational reproducibility in support of quality science is a long-standing goal in the field, but one where the very concept is confused. Computational workflow managers are among the most powerful tools available in this quest, which, alongside contemporary software engineering techniques and provenance capture, are moving towards a world of fully Rerunnable, Repeatable, Reproducible, and Replicable research. Claerbout [8] coined the concept of computational reproducibility in scientific computing, and Peng [9] elaborates on what is needed to achieve computational reproducibility in science. Drummond [10] separates the concepts of reproducibility and applicability. Barba [5] reviews the myriad of terminologies used to describe computational reproducibility as Benureau and Rougier [11] provide a framework separation Rerunning, Repeating, Reproducing, and Replicating computational works. Central to these discussions is a separation between bitwise precise replication of exact results and conceptual reproductions of an existing work independent of precise implementation. Our definitions of these Reproducibility tenets unify these two views. Current suggested best practices in pursuit of scientific reproducibility include simple rules [12], adapting modern tooling from industry including notebook environments [13, 14, 15], sharing code [16] and publishing openly [17, 18]. Related to the signature approach we present in this work, Bellini [19] investigates the feasibility of building a DOI-like dataset identifier leveraging public blockchains.
@@
Highly scalable and trustable computing is critical to conquering our largest scientific endeavors [26], and as we move towards reproducible computational workflow management, we must also ensure these efforts function as expected; a key motivation behind our signature-based approach to workflow reproducibility.
@@ 3 Computational Workflows
Computational workflows (herein, ‘workflows’) are structured assemblies of computational tasks and data stores [27] and underpin some of the world’s most extensive science projects, including the SKA [4]. Workflows separate domain-specific software knowledge from scientific methodology, making increasingly complex computations possible for increasingly more scientists, and have seen broad uptake in the past three decades [28]. Many workflow management systems provide several flavors of workflows for specific science domains [29], and thus, any attempt to address workflow reproducibility should be tool-agnostic. This section introduces a generic computational workflow model and discusses potential mappings to several well-known workflow management systems before delving into a concrete mapping for DALiuGE [7], [30], a workflow management system developed to execute workflows at the scale the SKA requires. Providing a general workflow model allows reasoning about computational reproducibility applied to the model to apply to other mapped workflow systems. Such interoperability of reasoning is necessary in working towards a holistic approach to computational reproducibility in data-intensive sciences.
@@ 3.1 Scale and System Agnostic Workflow Model
Figure 1 presents a generalized Unified Modeling Language (UML) workflow model. We intend for this model to make reasoning about testing workflow executions concrete and applicable to arbitrary workflow management systems. Introducing a universal representation makes discussing, understanding, and implementing subsequently described reproducibility tests easier.
Figure 1: A UML diagram depicting an arbitrary-scale workflow definition.
@@
A logical task describes a task characterized by high-level information like programming language, algorithm, and required options and parameters.
Logical task fields capture additional detail about a logical task, such as the exact software script, package, or command executed.
@@
Physical task fields capture additional detail about a physical task’s execution, such as the machine details a task is eventually executed on.
@@
A UML definition makes all modeled workflows well-defined software in their own right and permits native workflow implementation independent of using a workflow management system. Moreover, we separate logical and physical workflow components. This separation captures the intuitive experience of designing or describing workflows as a separate entity before executing a workflow and allows for reasoning and comparison of logical workflow designs irrespective of their eventual implementation in software as physical components. Moreover, several physical components can, in principle, implement a single logical workflow component, as is the case in large-scale workflows. Finally, the model’s requirement to alternate computing tasks and data artifacts ensures that a workflow description captures all interim and transient data artifacts. This lightweight model aims to flexibly model most existing workflow management system behaviors, capturing the difference between designing and executing workflows and between elements that compute and those that store data. While further model formalization, including ontologies for generalized scientific processes, allows accurate descriptions of data provenance [31], further formalization is beyond the scope of this research, whose focus lies on direct software implementation and testing of high-level reproducibility concerns.
3.1.1 Mapping the workflow model to workflow management systems
Table 1: Mapping from workflow model to workflow management system components. Although this work presents only an implementation for DALiuGE, we intend that the reasoning applied to the workflow model can be applied to other workflow management systems.
Workflow
Model
Element
DALiuGE Taverna Askalon Pegasus Kepler Nextflow REANA
Component Component - Action - Component - -
Logical
Task
Logical
Component
Activity
Activity
Type
Abstract
Task Nodes
Actor/Director Process
Container
Description
Physical
Task
Application
Drop
Processor
Activity
Deployment
Executable
Task Nodes
Executable Process
Container
Instance
Logical
Data Artifact
Logical
Data Component
Data Link Data Ports
Edges /
Logical Filenames
Data Ports Filename Filename
Physical
Data Artifact
Data Drop Data Item Systems Files System Files System Files
File /
Input Stream
File Contents
Logical
Workflow
Logical Graph
Workflow
Object (model)
Compound
Activity
Abstract
Workflow
Model Pipeline
Workflow
Description
Physical
Workflow
Physical Graph
Executable
Workflow Graph
Activity Set
Executable
Workflow
Workflow Workflow Workflow
Mapping between specific workflow management systems and a generic workflow model facilitates a consistent approach to reasoning about reproducibility. While reasoning about any possible workflow management systems is likely impossible, Table 1 presents possible mappings for several well-known workflow management systems. Aiming to provide a good cross-section of approaches and use cases for workflow management systems, the review da Silva et al. provide [23] motivates our selection. For each workflow management system presented, we provide a brief description of its history, usage, and reproducibility efforts concerning that particular system.
Taverna was among the first scientific workflow management systems designed primarily for orchestrating web services as data-processing workflows [23]. Its web-service-oriented approach was widely adopted in bioinformatics and influenced subsequent workflow management systems. Taverna’s ability to package workflows as separate research objects sets a precedent for workflow sharing and reuse. In mapping to the generic workflow model, Taverna workflows are expressed as graphs of computational components, typically web services, with a distinction between logical and physical data captured through a data manager [32]. Taverna’s early adoption of workflow provenance remains a core feature [33, 34], though it does not explicitly focus on reproducibility and reusability.
Askalon is a pioneering workflow management system focused on grid computing, utilizing AGWL, an XML schema for defining workflows [35]. Askalon separates workflow design from execution concerns, allowing novel scheduling and runtime environments. Its robust schema-based approach makes it an ideal candidate for mapping to a generic workflow model, particularly in expressing parallelism and scalability. Mapping Askalon to the model is straightforward due to its structured approach, though the lack of explicit logical data objects in place of data ports poses challenges. Despite its pre-reproducibility crisis design, Askalon’s rigorous workflow schema guarantees a level of correctness and uniqueness.
Pegasus is a workflow management system designed for executing workflows in distributed environments, with a strong focus on scalability and reliability [36, 37]. It has been widely used across various scientific domains, making it a canonical choice for analysis [23]. Pegasus workflows are expressed as directed acyclic graphs (DAGs) in XML, which aligns well with the generic workflow model’s structure. This alignment facilitates the mapping process, especially in distinguishing between logical and physical tasks. Pegasus’s intelligent features, such as the replica catalog [37], enhance performance and reduce recomputation, though the dynamic nature of distributed environments poses challenges for reproducibility.
Kepler is a scientific workflow management system that leverages a unique actor-director model inherited from the Ptolemy II Java library to support a wide range of scientific domains [38, 28]. The flexibility of this model allows Kepler to adapt workflows without altering their logic, making it a significant system for mapping to the generic workflow model. Kepler’s approach to workflow enactment, coupled with its forward-thinking data provenance system [39, 40, 41], ensures that workflows can be partially Rerun, enhancing reproducibility. However, the actor-director paradigm introduces complexity in mapping logical tasks to physical executions within the generic workflow model.
Nextflow is a modern workflow management system built from open-source components, particularly popular in bioinformatics for its performance and a wide array of supported executors [42]. It uses a domain-specific language to assemble workflows from processes and files, which aligns well with the generic workflow model. Nextflow’s emphasis on containerization and reproducibility makes it straightforward to map its elements to the model, ensuring that workflows are both scalable and reproducible. The success of repositories like nf-co.re [43] illustrates the practical benefits of integrating reproducibility directly into workflow management systems.
REANA is a newer workflow management system designed with a focus on reproducibility in data-intensive sciences, particularly high-energy particle physics [44, 45]. Built on modern open-source tools like Kubernetes and Docker, REANA represents the state-of-the-art in workflow management. Mapping REANA to the generic workflow model is facilitated by its use of directed acyclic graph (DAG) models like CWL and Yadage, making the mapping process relatively straightforward. However, REANA’s simplicity, while robust, limits its flexibility in separating logical and physical workflow components, making reproducibility comparisons a manual task.
3.1.2 Concrete mapping to DALiuGE
@@
DALiuGE [7], [30] supports workflow management at SKA scale. The unique data-driven design, decentralized execution, and careful share-nothing architecture enable this. Moreover, DALiuGE strictly separates logical workflow design from physical workflow execution and thus allows astronomers to design workflows without intimate component knowledge and, more importantly, facilitates the broad reuse of logical workflows. For this study, we have implemented the generalized ideas described in this work in DALiuGE since we have complete code control and in-depth knowledge of the code structure. We, therefore, further elaborate on the workings of DALiuGE and its mapping to the generic workflow model. DALiuGE specifies workflows as graphs of components and separates workflows into two main layers, logical and physical, and then sub-divides each of those into template and non-template versions. Workflows are developed as logical templates comprised of exclusively logical components and control structures. Scientists will design a Logical Graph Template (LGT) several months or years before the actual processing and for extensive reuse. LGTs expose configuration options to make them more specific for concrete execution. A Logical Graph (LG) is a LGT with a concrete configuration. LG specification occurs on the order of weeks before execution. DALiuGE then unrolls the LG into a Physical Graph Template (PGT), resolving control structures and presenting the net computation required. This process is unique for each LG, but the result can vary depending on the translator algorithm used. A physical graph template is a directed acyclic graph (DAG), a common representation for computational workflows [23]. During the scheduling process, DALiuGE partitions a PGT into sub-graphs based on available compute resources, resulting in a Physical Graph (PG). This step is performed just before execution since it maps the physical workflow components to the available resources. Each component of the DAG is a fully specified computational task or data artifact. Finally, the DALiuGE daemon triggers the first physical components in a graph, but after, components communicate with each other in a fully decentralized manner. Components can store runtime-specific information such as log files or data summaries, which together form a Runtime Graph (RG) after graph execution has been completed.
The first column of Table 1 summarizes the mapping of the generic workflow model to DALiuGE components. This mapping makes implementing reproducibility testing mechanisms feasible, and the already present distinction between ahead-of-time ‘logical’ workflow design and runtime ‘physical’ workflow execution makes mapping and subsequent implementation straightforward. However, including a generic model and candidate mappings to other workflow management systems encourages further work.
@@ 4 Reproducible Workflows
Contemporary science is complex; variability in the design and execution of complex data analysis pipelines can be scientifically significant [46]. Most attempts to ensure workflows are reproducible involve analyzing artifact provenance, describing precisely how pieces of data came to exist [31]. We build on this lineage, but instead of designing provenance capture to describe the entire technical process, we formulate reproducibility tests, singling out parts of workflow executions expected to be kept constant between executions. Doing so allows for flexible workflow characterization and does not limit this testing-based approach to a specific workflow management system, a critical conceptual requirement for the multi-decade-spanning SKA project and the broader scientific field. Exploiting the structured nature of workflows provides constant time provenance comparison, which is suitably efficient for deployment at extreme data scales.
@@ 4.1 Reproducibility Tenets
Melding computing into satisfyingly reproducible science is a well-discussed and unsolved problem [1, 47, 6]. Current efforts range from asserting standards on individual pieces of code [11], enacting well-defined data management protocols [22, 48], containerizing whole workflows [45], measures on behalf of journal publishers [49], and conference organizers [50] and careful provenance production and collection [31]. While certainly practical, we propose that forward-facing measures insufficiently integrate computing into science; they are ultimately empirical and require an inductive step to evaluate their efficacy. The proposed approach demarcates implementation from abstract computational ideas, data and method reproduction, workflow similarities between workflow management systems, and any combination of these traits. We present seven reproducibility tenets formulated as workflow component invariants, each defining a testable assertion rather than a series of prescribed rules. For each tenet, we provide an unambiguous definition, an explanatory UML diagram based on the generic workflow model, a justification, and a few motivating example scenarios, providing practical motivation behind our definition.
@@
We use the popular terminology to Rerun, Repeat, Recompute, Reproduce, and Replicate computational workflows as these terms capture intuitive concepts around scientific reproducibility [51, 11, 5, 44]. Table 2 summarizes the definitions, and the following subsections provide detailed explanations of each tenet in addition to the formal UML diagram for each of them. The diagrams are structurally identical to Figure 1, with the items required to be kept identical to satisfy that particular tenet highlighted in blue. Later, Section 5.2 describes the precise provenance information included in a concrete mapping to DALiuGE to satisfy each tenet.